the length of a rectangle is 5cm longer than its width the area is 66 cm square find the perimeter of this rectangle
Answer by lwsshak3(11628)

the length of a rectangle is 5cm longer than its width the area is 66 cm square find the perimeter of this rectangle
***
let x=width of rectangle
x+5=length of rectangle
length*width=area
x(x+5)=66
x^2+5x-66=0
(x+11)(x-6)=0
x=6
x+5=11
width of rectangle=6
length of rectangle=11
2*(width+length)=perimeter
2(6+11)=34
perimeter of rectangle=34 cm